# Mosaic-IT: Free Compositional Data Augmentation Improves Instruction Tuning

### v2

@@ 4.2 Ablation studies
In this section, extensive ablation experiments are conducted on Mistral-7B using with the Alpaca-GPT4 dataset to verify our method. We utilize Pair-wise comparison for evaluation.
(a) Ablation on Mosaic-IT strategies.
@@
(b) Ablation on the Max Number of Instructions.
@@
Ablation on Mosaic Strategies is presented in Table 4(a). Primary represents the Primary Mosaic Strategy. The winning score of this setting is greater than $1.0$, indicating a better performance compared with the baseline method. This comparison directly verifies the effectiveness of the idea of introducing multiple instructions during training, which complicates the instructions at no cost and improves the instruction-following ability of LLMs. Format represents the Format Strategy. Although the winning score is only slightly greater than the naive version, this version makes it possible for LLMs to follow the customized user-defined formats, indicating great potential for the controllability of LLMs. Moreover, the format version can be easily used with other types of meta instructions, showing great extensibility. Permute represents the Permute Strategy that builds on the Format Strategy with a probability of $\nicefrac{{1}}{{2}}$, similar to Maskout. Permute/Maskout represents our default setting, where the Permute or Maskout Strategies are utilized together with the Format Strategie with a probability of $\nicefrac{{1}}{{3}}$. All these $3$ settings show higher performance than the format version, indicating the effectiveness of Advanced Mosaic Strategies which define more complicated meta instructions.
Table 5: Ablation on the Distribution of Number of Instructions. The distribution formula and data counts for different settings are shown in Appendix A. “Mix $\leq 5$” represents the percentage of samples with the number of instructions less or equal to $5$.
@@
Ablation on the Max Number of Instructions is presented in Table 4(b), including the pair-wise comparison values. As shown in the table, when the max number is set as $2$, i.e. at most $2$ instructions/responses are concatenated together, the performance is almost the same as the baseline, indicating the ineffectiveness. However, when the max number grows, the corresponding winning scores also grow consistently. This trend shows that the more instructions concatenated together, the better the instruction-following ability. We hypothesize that, with the growth of the number of instructions, the overall instruction becomes much harder to follow, especially for the permute and maskout strategies, which benefits LLMs’ instruction-following capability.
Figure 3: Ablation on the Distribution of Number of Instructions, the visualization of distribution comparisons.
@@
Ablation on the Distribution of Number of Instructions is presented, including the pair-wise comparison values in Table 5 and detailed number distribution comparisons in Figure 3, which aims at identifying how this count distribution affects the performance of our method. The detailed distribution formula and data counts are provided in the Appendix A. Fix represents the setting where all the overall instructions are concatenated with a fixed number of instructions, which we set as $10$ unless the overall instructions exceed the max length limit. Exponential represents the setting where the number of instructions is sampled following the exponential distribution. Under these two settings, less than $3\%$ of the overall instructions are concatenated by less or equal to $5$ original instructions. The lack of few-instruction concatenated samples negatively affects the LLMs’ ability to follow the single instruction, which is employed by most of the existing evaluation methods, leading to worse performances. Pareto, Log-normal, and Logistic represents the corresponding distribution that are utilized for sampling. Different from the above two settings, approximately $10\%$ of the overall instructions are composed of fewer original instructions, thus ensuring the LLMs are trained with samples with sufficiently diverse lengths, resulting in optimal performances. Uniform is our default setting, representing using the uniform distribution where different numbers are sampled evenly. In this situation, the LLMs are trained with samples with the most diverse lengths, thus avoiding the LLMs overfit to simple lengthy responses.
5 Further Discussion
5.1 Preliminary Experiments: Performance Degradation
GPT-3.5-turbo (Sequential) 0.357 0.014 0.336 0.055 0.303 0.064
GPT-3.5-turbo (Random) 0.315 0.124 0.330 0.156 0.198 0.312
GPT-4-turbo (Sequential) 0.176 0.000 0.137 0.000 0.140 0.000
GPT-4-turbo (Random) 0.139 0.000 0.153 0.014 0.101 0.005
5.2 Mosaic: Alleviating Memorizing
In the original instruction tuning process, each data sample will be trained several times for LLMs without changes to the instructions and responses. This training process poses risks to the potential memorizing effects on training samples, which can be partially indicated by the “stair-like” training loss curves as shown in Figure 4. In the figure, all the training settings are kept the same between the baseline models and Mosaic-IT models, including the Learning Rate, Warm-up Ratio, Learning Rate Schedule (Cosine), Batch Size, etc. For the baseline methods, the training loss hardly decreases within each epoch of training but drops dramatically when the LLMs meet the same training samples again, which indicates a potential memorizing effect of training samples and potential overfitting. However, when utilizing our method, the random mosaics of original instructions with diverse and complex meta-instructions largely diversify the overall training instructions. Although each original data sample will still be seen by LLMs several times during training, the overall context varies dramatically as each original sample is only an atomic element of the overall mosaic sample, indicating that there will be no identical overall instructions during the whole training process. Thus this augmentation largely alleviates the potential memorizing and overfitting problems as shown in the figure, where the training loss decreases smoothly, representing the gradual learning process.
5.3 Mosaic: Improving Efficiency
Time (min) 827 121 129 133 133 143 716 426 305 245 202 173
One of the benefits of our method is the efficiency of the training process. Given an existing dataset, our mosaic processes largely decrease the number of total overall instructions and the total number of gradient descents, leading to a reduction in the training process. The detailed comparison is shown in Table 7, which is based on the Mistral-7B model on the Alpaca-GPT4 dataset. The time is calculated based on four NVIDIA A100 Graphic Cards. As shown, our method greatly decreases the training time to approximately $16\%$ to $25\%$ while achieving better performances, especially when there are mosaic samples with larger permutation counts.
5.4 Mosaic: Why It Works?
The effect of our method is aligned with the Dense and Aligned Captions Doveh et al. (2023) used in VL Models, which utilizes denser captions to promote the VL models. Different from all previous methods which require LLMs to generate the whole response conditioned on the whole instruction, our method forces LLMs to generate responses selectively conditioned on some critical parts of the overall dense instruction. Especially in advanced strategies, LLMs are required to generate responses conditioned not only on sequential parts of the instruction, but diverse and randomized segments of it, which is defined by meta-instruction.
Compared to original settings, our setting requires LLMs to first understand the formats and orders defined in the meta-instructions, then adapt to conditioning on different parts of the instructions when generating responses. This process forces LLMs to develop a more comprehensive understanding of context, prioritize various pieces of information, and manage complex dependencies between instructions, thus improving instruction-following performance. Moreover, the entire process is data- and model-agnostic, ensuring the generalizability of our method.
5.5 Limitation and Future Directions
6 Conclusion
We introduce Mosaic Instruction Tuning (Mosaic-IT), a novel, human/model-free method to enhance instruction tuning for LLMs. By concatenating multiple instruction-response samples and using higher-level meta-instructions, Mosaic-IT improves multi-step and format-following capabilities. Our evaluations show superior performance and an $80\%$ reduction in training costs compared to the original methods. Mosaic-IT’s simplicity and efficiency make it a scalable solution for improving LLMs without extensive human intervention or resource-intensive teacher models. Our results highlight the potential of innovative data augmentation techniques in advancing LLM capabilities.
Appendix A Detailed Distribution for Ablation on Mixture Distribution
A.1 Distribution description
@@
where $\lambda=1$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater then $k_{max}$.
@@
Pareto Distribution [^4] https://numpy.org/doc/stable/reference/random/generated/numpy.random.pareto.html: The Pareto II or Lomax distribution is a shifted Pareto distribution. It can be considered as a simplified version of the Generalized Pareto distribution, with the scale set to one and the location set to zero. The probability density function (PDF) for the Pareto distribution is:
@@ A.2 Distribution visualization
Appendix B Experimental Setup
B.1 Implementation Details
For the three base pre-trained models, Llama2-7B, Llama2-13B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023), we utilize the prompt and code base from Vicuna Chiang et al. (2023) and flash attention Dao et al. (2022). The overall training arguments are aligned with the common training configuration. The Adam optimizer Kingma & Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. When training the baseline models Llama2-7B and Llama2-13B, the maximum learning rate is set to $2\times 10^{-5}$ with the warmup rate as $0.03$ for $3$ epochs. When training the baseline models on Mistral-7B, the maximum learning rate is set to $1\times 10^{-5}$ with the warmup rate as $0.1$ for $3$ epochs. For the three models, Llama-3-8B Dubey et al. (2024), Phi-3 Abdin et al. (2024), and Gemma2-2B Team et al. (2024), we utilize the code base from LLaMA-Factory Zheng et al. (2024b). The max token length is set with $4096$ following the modern settings and we train the model for $2$ epochs. Other parameters are kept the same as the above.
B.2 Training Dataset
The Alpaca dataset Taori et al. (2023) comprises $52,000$ instruction-following samples and is constructed utilizing the self-instruct paradigm Wang et al. (2023b). This dataset was produced by employing OpenAI’s text-davinci-003 model. Characterized as a classical dataset with moderate quality attributes, the Alpaca dataset serves as an initial platform to validate our methodology. To further substantiate our approach using a dataset of inherently high quality, we also applied our method to the Alpaca-GPT4 dataset Peng et al. (2023), which features responses generated by GPT4. The WizardLM dataset Xu et al. (2023) is also utilized in our method, which contains $70,000$ samples created by the evolution algorithm proposed by them. With ChatGPT-3.5 utilized, the data quality on WizardLM is largely guaranteed. The Vicuna 1M dataset Zheng et al. (2024a) is a large-scale dataset containing one million real-world conversations with 25 state-of-the-art LLMs, due to the computation budget, $300$k instances are randomly sampled for our experiments. Magpie dataset Xu et al. (2024b) is a most recent SOTA synthetic dataset with $300$k samples.
B.3 Evaluation Metrics
Pair-wise Comparison by using powerful LLMs like GPT-4 is recently widely accepted and becoming a common practice Touvron et al. (2023b); Chiang et al. (2023); Dettmers et al. (2023); Liu et al. (2023b); Chiang & Lee (2023). The evaluation of responses from LLMs, especially in open-domain contexts where definitive ground truth is hard to establish, continues to be an intricate and evolving research domain. Recent studies, however, have indicated a notable alignment between GPT-4’s performance evaluations and human assessments Zheng et al. (2023); Li et al. (2023c); Sottana et al. (2023), thereby establishing a credible foundation for this evaluative methodology. We adopted test instruction sets from WizardLM Xu et al. (2023), comprising 218 diverse, human-curated instructions for pair-wise comparison. We directly follow the evaluation framework proposed by Chen et al. (2023); Li et al. (2024d), which evaluates responses on a scale spanning from 1 to 10 across multiple dimensions. To further address positional bias, as discussed by Ko et al. (2020); Wang et al. (2023a), the comparison is conducted in two distinct sequences, LLM1’s response first and then LLM2’s response first, ensuring a fair assessment of model performance. Evaluation outcomes are categorized into ’win-tie-loss’ for each instruction. The detailed evaluation prompt is provided in Appendix E.
Appendix C Related Work
@@
Instruction Data Improvement: WizardLM Xu et al. (2023) first proposes an Evol Algorithm to complicate the existing data and reach supreme performance. LaMini-LM Wu et al. (2024) innovatively generates ”Topic-Guided” instructions utilizing Wiki data. Tree-Instruct Zhao et al. (2024) preliminarily explores the relationship between instruction complexity and Alignment and proposes adding nodes to complicate the instruction. UltraChat Ding et al. (2023) establishes broad thematic scopes, systematically generating numerous instructions within each. Reflection-Tuning Li et al. (2023a) sequentially refines both instructions and responses by focusing on specific evaluative criteria. DEITA Liu et al. (2023a) utilizes ChatGPT to diversify and then select the data. Selective Reflection-Tuning Li et al. (2024b) proposes a teacher-student collaborative pipeline to improve and select the data. Instruction Fusion Guo et al. (2024) proposes to utilize ChatGPT4 to merge two distinct instructions for further complexity enhancement. These advancements showcase a shift towards automating the generation and refinement of datasets, reducing reliance on human labor.
Instruction Data Selection: It is widely accepted that ”quality is all you need” Touvron et al. (2023b); Zhou et al. (2023) for instruction tuning. LIMA Zhou et al. (2023) demonstrates that merely 1,000 human-carefully-curated, high-quality training instances can substantially enhance the instruction-following performance. InsTag Lu et al. (2023) employs the proprietary model, ChatGPT, to tag instruction data and select data with complex tags. Alpagasus Chen et al. (2023) utilizes proprietary LLMs chatGPT and Claude2 to directly assess the quality of instruction tuning data. Cherry LLM Li et al. (2024d) proposes the Instruction-Following Difficulty (IFD) scores to assess the difficulty of the instructions, which is a self-guided method in which no extra LLMs are utilized. Motivated by Humpback Li et al. (2023b), Selective Reflection-Tuning Li et al. (2024b) extends the IFD score to a reverse version, focusing on the feasibility of responses. Du et al. (2023) and Bukharin & Zhao (2023) utilize reward models as the base scores for measuring data quality. DEITA Liu et al. (2023a) experiments on several different data selection metrics and builds a dataset with high quality. Superfiltering Li et al. (2024c) reveals the consistency between weak and strong language models in perceiving instruction difficulty, making the filtering process much more efficient. All these works are devoted to distinguishing and selecting good data samples from bad ones for instruction tuning.
Appendix D Predefined Rules
@@
Appendix E Prompt for Evaluation
@@
Appendix F Detailed Performance Scores on Llama3, Phi3 and Gemma2

### v3

@@ Appendix C Detailed Ablation Studies
In this section, extensive ablation experiments are conducted on Mistral-7B using the Alpaca-GPT4 dataset to verify our method. We utilize Pair-wise comparison for evaluation.
C.1 Ablation on Mosaic Strategies
Ablation on Mosaic Strategies is presented in Table 6(a). Primary represents the Primary Mosaic Strategy. The winning score of this setting is greater than $1.0$, indicating a better performance compared with the baseline method. This comparison directly verifies the effectiveness of the idea of introducing multiple instructions during training, which complicates the instructions at no cost and improves the instruction-following ability of LLMs. Format represents the Format Strategy. Although the winning score is only slightly greater than the naive version, this version makes it possible for LLMs to follow the customized user-defined formats, indicating great potential for the controllability of LLMs. Moreover, the format version can be easily used with other types of meta instructions, showing great extensibility. Permute represents the Permute Strategy that builds on the Format Strategy with a probability of $\nicefrac{{1}}{{2}}$, similar to Maskout. Permute/Maskout represents our default setting, where the Permute or Maskout Strategies are utilized together with the Format Strategy with a probability of $\nicefrac{{1}}{{3}}$. All these $3$ settings show higher performance than the format version, indicating the effectiveness of Advanced Mosaic Strategies, which define more complicated meta instructions.
C.2 Ablation on the Max Number of Instructions
Ablation on the Max Number of Instructions is presented in Table 6(b), including the pair-wise comparison values. As shown in the table, when the max number is set as $2$, i.e., at most $2$ instructions/responses are concatenated together, the performance is almost the same as the baseline, indicating the ineffectiveness. However, when the max number grows, the corresponding winning scores also grow consistently. This trend shows that the more instructions concatenated together, the better the instruction-following ability. We hypothesize that, with the growth of the number of instructions, the overall instruction becomes much harder to follow, especially for the permute and maskout strategies, which benefit LLMs’ instruction-following capability.
@@ C.2 Ablation on the Max Number of Instructions
(a) Ablation on Mosaic-IT strategies.
@@
(b) Ablation on the Max Number of Instructions.
@@
Table 7: Ablation on the Distribution of Number of Instructions. The distribution formula and data counts for different settings are shown in Appendix E. “Mix $\leq 5$” represents the percentage of samples with the number of instructions less or equal to $5$.
Figure 6: Ablation on the Distribution of Number of Instructions, the visualization of distributions.
C.3 Ablation on the Distribution of Number of Instructions
@@
Ablation on the Distribution of Number of Instructions is presented, including the pair-wise comparison values in Table 7 and detailed number distribution comparisons in Figure 6, which aims at identifying how this count distribution affects the performance of our method. The detailed distribution formula and data counts are provided in the Appendix E. Fix represents the setting where all the overall instructions are concatenated with a fixed number of instructions, which we set as $10$ unless the overall instructions exceed the max length limit. Exponential represents the setting where the number of instructions is sampled following the exponential distribution. Under these two settings, less than $3\%$ of the overall instructions are concatenated by less than or equal to $5$ original instructions. The lack of few-instruction concatenated samples negatively affects the LLMs’ ability to follow the single instruction, which is employed by most of the existing evaluation methods, leading to worse performance. Pareto, Log-normal, and Logistic represents the corresponding distribution that are utilized for sampling. Different from the above two settings, approximately $10\%$ of the overall instructions are composed of fewer original instructions, thus ensuring the LLMs are trained with samples with sufficiently diverse lengths, resulting in optimal performances. Uniform is our default setting, representing using the uniform distribution where different numbers are sampled evenly. In this situation, the LLMs are trained with samples with the most diverse lengths, thus avoiding the LLMs overfitting to simple lengthy responses.
Method Alpaca Eval 2 (LC) Alpaca Eval 2 Pair-wise Compare (with non-mosaic) Pair-wise Compare (with pure-random)
Pure-random Concatenation 5.00 7.81 1.349 1.000
Concatenation with Semantic Groups 7.80 6.51 1.275 0.936
Table 8: Comparison with Semantic grouping
C.4 Ablation on Semantic Grouping
To demonstrate that including other concatenation methods could further strengthen our argument, we further conduct experiments using a semantic grouping approach based on the Alpaca-GPT4 dataset to fine-tune Mistral.
Specifically, we utilize a sentence transformer model all-mpnet-base-v2111https://huggingface.co/sentence-transformers/all-mpnet-base-v2 to obtain the semantic embedding for each sample in the dataset, and then we apply the K-means algorithm to group these data samples into multiple clusters. To ensure enough samples per cluster, we set K=52 as the dataset contains 52k samples in total. Given the clusters, each concatenated sample is composed of multiple samples randomly drawn from the same cluster. We keep using the same training hyperparameters as in our main experiments. As shown in Table 8, we report the performance on two evaluation metrics: pair-wise comparison and Alpaca Eval.
The semantic concatenation can still outperform the non-mosaic baseline by a large margin, indicating the effectiveness and potential of our Mosaic-IT augmentations and tasks. The semantic concatenation method has a slightly lower performance than the pure-random concatenation method, on pair-wise comparison and Alpaca Eval 2 scores. However, it achieves a much higher Alpaca Eval 2 (LC) score. This result suggests that the response quality of the model trained with semantic concatenation is on par with pure-random, but the response length is shorter and more condensed. We found that semantic grouping leads to clusters with highly different average lengths of samples: The largest average length is 316.7 tokens, while the smallest is 31.4 tokens. This discrepancy makes the lengths of Mosaic-IT concatenated samples more diverse, resulting in a better trade-off between the quality and length of the responses.
Appendix D Multi-Instruction Evaluation
To verify our trained models’ capability to follow multiple instructions and meta-instructions in one inference, we create a test set of compositional instructions from WizardLM test sets using Mosaic-IT. For simplicity, we name this new test setting as Mosaic Task, which evaluates LLMs’ capability to follow multiple instructions with additional diverse constraints (meta-instructions). One example of Mosaic Task is shown as follows.
  Example of Mosaic Task
  System Prompt
You are a helpful and precise assistant for providing the answer.
User Prompt
Respond to each of the following instructions in reverse of the original order.
[Ins1]
[Ins2]
[Ins3]
We use the success rate(%) to evaluate the performance of models on the Mosaic task. A response is successful if it follows the meta-instruction and no instruction is ignored (unless the meta-instruction masks it). In the table below, we report the success rate (%) of LLMs following three meta-instruction strategies, i.e., Format, Permute, and Maskout, on compositional augmentations of different numbers of instructions (i.e., 3, 5, 7 instructions). We report the success rates of GPT4o, two base models, and their Mosaic-IT finetuned versions, as shown in Table 9.
The results expose the weaknesses of existing LLMs on Mosaic-IT tasks and show that training on Mosaic-IT augmentations can significantly improve performance. Specifically, existing LLMs, even GPT4o, can not perfectly follow multiple instructions with diverse constraints, not to mention other open-source models like Llama3 finetuned on datasets such as Magpie. These results further demonstrate the difficulty and complexity of Mosaic-IT tasks for existing LLMs, indicating the novelty of our method. The compositional reasoning capability required by Mosaic-IT tasks cannot be covered by the capabilities of base LLMs and existing instruction-tuning datasets. For example, the success rates of Mistral + Alpaca-GPT4 (baseline) and Llama3 + Magepie (baseline) are similar, although Llama3 + Magepie has relatively better general instruction-following capabilities among them.
Our method can bridge the significant gap and enhance LLMs’ capability to follow multiple instructions with diverse constraints. Moreover, our data augmentation is cost-free and does not take any effort from humans or models.
Model 3 Instructions 5 Instructions 7 Instructions
Format Permute Maskout Format Permute Maskout Format Permute Maskout
GPT4o 59.17 55.05 41.46 56.88 51.38 26.13 29.82 37.16 24.27
Mistral + Alpaca-GPT4 (baseline) 20.18 3.67 3.25 10.09 2.75 5.41 7.34 0.92 0.97
Mistral + Alpaca-GPT4 (mosaic) 98.32 66.51 69.11 95.87 60.55 67.57 97.25 64.68 66.02
Llama3 + Magepie (baseline) 16.06 8.26 7.32 9.63 1.38 5.41 5.50 2.75 3.88
Llama3 + Magepie (mosaic) 97.71 79.82 84.55 94.95 72.94 77.48 76.61 61.01 85.44
Table 9: Performance comparison across multiple instruction settings.
Appendix E Detailed Distribution for Ablation on Mixture Distribution
E.1 Distribution description
@@
where $\lambda=1$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater than $k_{max}$.
@@
Pareto Distribution [^5] https://numpy.org/doc/stable/reference/random/generated/numpy.random.pareto.html: The Pareto II or Lomax distribution is a shifted Pareto distribution. It can be considered a simplified version of the Generalized Pareto distribution, with the scale set to one and the location set to zero. The probability density function (PDF) for the Pareto distribution is:
@@ E.2 Distribution visualization
Appendix F Related Work
@@
Instruction Data Improvement: WizardLM Xu et al. (2023) first proposes an Evol Algorithm to complicate the existing data and reach supreme performance. LaMini-LM Wu et al. (2024) innovatively generates "Topic-Guided" instructions utilizing Wiki data. Tree-Instruct Zhao et al. (2024) preliminarily explores the relationship between instruction complexity and Alignment and proposes adding nodes to complicate the instruction. UltraChat Ding et al. (2023) establishes broad thematic scopes, systematically generating numerous instructions within each. Reflection-Tuning Li et al. (2023a) sequentially refines both instructions and responses by focusing on specific evaluative criteria. DEITA Liu et al. (2023a) utilizes ChatGPT to diversify and then select the data. Selective Reflection-Tuning Li et al. (2024c) proposes a teacher-student collaborative pipeline to improve and select the data. Instruction Fusion Guo et al. (2024) proposes to utilize ChatGPT4 to merge two distinct instructions for further complexity enhancement. These advancements showcase a shift towards automating the generation and refinement of datasets, reducing reliance on human labor.
Instruction Data Selection: It is widely accepted that "quality is all you need" Touvron et al. (2023b); Zhou et al. (2023) for instruction tuning. LIMA Zhou et al. (2023) demonstrates that merely 1,000 human-carefully-curated, high-quality training instances can substantially enhance the instruction-following performance. InsTag Lu et al. (2023) employs the proprietary model, ChatGPT, to tag instruction data and select data with complex tags. Alpagasus Chen et al. (2023) utilizes proprietary LLMs chatGPT and Claude2 to directly assess the quality of instruction tuning data. Cherry LLM Li et al. (2024e) proposes the Instruction-Following Difficulty (IFD) scores to assess the difficulty of the instructions, which is a self-guided method in which no extra LLMs are utilized. Motivated by Humpback Li et al. (2023b), Selective Reflection-Tuning Li et al. (2024c) extends the IFD score to a reverse version, focusing on the feasibility of responses. Du et al. (2023) and Bukharin and Zhao (2023) utilize reward models as the base scores for measuring data quality. DEITA Liu et al. (2023a) experiments on several different data selection metrics and builds a dataset with high quality. Superfiltering Li et al. (2024d) reveals the consistency between weak and strong language models in perceiving instruction difficulty, making the filtering process much more efficient. A recent work Li et al. (2025) tries to unify the different data evaluation metrics through the lens of gradients. All these works are devoted to distinguishing and selecting good data samples from bad ones for instruction tuning.
Appendix G Predefined Rules
@@
Appendix H Prompt for Evaluation
@@
Appendix I Detailed Performance Scores on Llama3, Phi3 and Gemma2